OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. If x represents Sarah's age now, which of the following expressions represents Ralph's age in four years?

TheSmartOne (thesmartone):

Ralph is R Sara is S

TheSmartOne (thesmartone):

R=3S is the first equation.

OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. If x represents Sarah's age now, which of the following expressions represents Ralph's age in four years? 3x 6x 2x + 4 3x + 4

TheSmartOne (thesmartone):

R+4=2(S+4) is the second equation

TheSmartOne (thesmartone):

Remember that R and S is their current age.

TheSmartOne (thesmartone):

so let us distribute the second equation it will become R+4=2S+8

TheSmartOne (thesmartone):

Now lets plug in the first equation in to the second one

TheSmartOne (thesmartone):

3S+4=2S+8 S=4

TheSmartOne (thesmartone):

R=3S R=3(4) R=12

TheSmartOne (thesmartone):

It wants Ralphs age in 4 years so R+4 which is 12+4=16

TheSmartOne (thesmartone):

So for your answer choices plug in x=4 because that is Sara's age and see which one gives you 16.

TheSmartOne (thesmartone):

It should be 3(4)+4=16 So the answer is 3x+4